# Beyond Multi-Token Prediction: Pretraining LLMs with Future Summaries

### v1

@@ 5 Related Work
Pitfalls of Next-Token Prediction (NTP).
While next-token prediction (NTP) is the standard loss for modern LLMs, its shortcomings are increasingly evident. The main issue is the mismatch between training (teacher forcing) and inference (autoregression): during training, the model sees ground-truth tokens, which encourages learning spurious correlations or “shortcuts” rather than the true data distribution. Bachmann and Nagarajan (2024) describe this as the “Clever Hans cheat”, where the model exploits trivial cues from the prefix and fails on lookahead tasks such as their path-star graph problem. Nagarajan et al. (2025) further show that NTP is data-inefficient for tasks requiring a “leap of thought,” such as Sibling Discovery. Our approach tackles these failures by providing a more robust, long-range training signal that discourages such shortcuts.
Going Beyond Immediate Next Token.
@@
Multi-Token Prediction (MTP) addresses limitations of Next Token Prediction (NTP) by using auxiliary heads to predict future tokens (Gloeckle et al., 2024). DeepSeek-V3 (Liu et al., 2024) modifies these auxiliary heads to allow slight teacher forcing, but the key issue with both the approaches is that scaling the auxiliary heads for long-range dependencies is impractical. Recent work aims to improve efficiency and capture longer dependencies. Gerontopoulos et al. (2025) introduce register tokens, predicting tokens $k$ steps ahead without architectural changes. Thankaraj et al. (2025) insert lookahead tokens containing future subsequences, while Liu et al. (2025) use a leap-based strategy to predict non-sequential future tokens. However, heuristically sampling random future tokens still poses the risk of missing informative long-range signals. Our FSP-RevLM addresses this by predicting a learned summary of the future, such that it can extract meaningful long-range information.
Bidirectional models.
@@
Leveraging reverse, or right-to-left, order during training has shown empirical benefits across multiple learning paradigms. The Belief State Transformer (BST) (Hu et al., 2024) employs dual forward and backward encoders to predict both the next token after a prefix and the previous token before a suffix. This bidirectional training encourages the model to form a compact belief state, though it does not explicitly address teacher forcing. In contrast, our FSP-RevLM incorporates a reverse model with a different objective: mitigating dependence on teacher forcing. While trained on standard left-to-right sequences, FSP-RevLM aligns the forward model’s embeddings with those of a reverse model, effectively distilling the reverse order into the forward language model. Another related approach, Meet-in-the-Middle (MiM) (Nguyen et al., 2023), jointly trains forward and backward models with shared parameters and employs an agreement regularizer to align their output distributions. Our method differs in two key aspects: (1) we perform distillation with the reverse model rather than parameter sharing, and (2) we avoid the somewhat impractical assumption required by MiM that the forward and reverse output distributions must match exactly.
@@ 6 Conclusion
In this work, we highlighted a key limitation of existing multi-token prediction methods: the difficulty of scaling auxiliary heads for long-horizon future prediction. Towards this, we proposed Future Summary Prediction (FSP), a novel pretraining framework that shifts the auxiliary objective from predicting specific future tokens to predicting a (learned) summary of the future. Experiments on 8B models, using both hand-crafted and learned summaries, demonstrate that FSP delivers a stronger, more robust training signal, improving over NTP and MTP on challenging reasoning and coding tasks. Our findings indicate that focusing on abstract, predictable aspects of the future is a promising strategy for designing more efficient and effective pretraining objectives for next-generation large language models.
@@ Acknowledgments
The authors thank Sachin Mehta for their suggestions regarding the real-world pretraining experimental setup. The authors also gratefully acknowledge the helpful discussions with Mathurin Videau, Sharut Gupta, Sarthak Mittal, and Andrei Nicolicioiu.

### v2

@@ 5 Related Work
Beyond Multi-token Prediction. Multi-token Prediction (MTP) extends Next-token Prediction (NTP) by introducing auxiliary heads to predict future tokens (Gloeckle et al., 2024), thereby providing a richer learning signal. Variants such as DeepSeek-MTP (Liu et al., 2024), Joint-token Prediction (Ahn et al., 2025), and Next-latent Prediction (Teoh et al., 2025) differ in how they structure these auxiliary predictions, but they retain a common limitation: scaling auxiliary heads to capture long-range dependencies becomes impractical. More recent approaches attempt to improve efficiency by predicting randomly selected or non-sequential future tokens (Gerontopoulos et al., 2025; Thankaraj et al., 2025; Liu et al., 2025). However, heuristically sampling future tokens risks missing the most informative long-range signals. Our approach (FSP-RevLM) addresses this by predicting a learned future summary, such that it can extract meaningful long-range information.
Related to the proposed hand-crafted summary approach (FSP-BoW), Yin et al. (2024) and Frydenlund (2025) incorporate objectives that predict an unordered set-level summary of future tokens, conceptually similar to our FSP-BoW and its binary cross-entropy formulation. However, our formulation is more general as we allow re-weighting schemes such as tf-idf scores within the binary cross-entropy objective and is implemented via auxiliary heads to be consistent with the multi-token prediction architecture (Gloeckle et al., 2024). More importantly, our contribution is not limited to a specific BoW variant: we introduce a unified framework for future summary prediction that clarifies the trade-offs between different forms of future supervision. For instance, a BoW summary directly resolves the path–star graph failure by aggregating the long-range information, yet it remains suboptimal when only part of the future is relevant, as observed in our sibling discovery experiments. These insights motivate moving beyond hand-crafted summaries toward a learned future summary via reverse LM, which adaptively captures the most informative aspects of the future.
Finally, regarding SemFormer (Yin et al., 2024), beyond its BoW-style supervision, its central idea is to introduce special planning tokens that are trained to predict a latent embedding of the future sequence, encouraging planning. While conceptually related, it applies future supervision only at designated planning tokens and relies on auto-encoding objective to learn future embeddings. In contrast, our approach requires no special planning tokens, enforces future-summary alignment at every position in the sequence through auxiliary heads, and learns future embeddings via the ReverseLM, yielding a fundamentally different supervision signal.
Leveraging the right-to-left signal.
@@
Leveraging reverse, or "right-to-left", order during training has shown empirical benefits across multiple learning paradigms. The Belief State Transformer (BST) (Hu et al., 2024) employs dual forward and backward encoders to predict both the next token after a prefix and the previous token before a suffix. This bidirectional training encourages the model to form a compact belief state, though it does not explicitly address teacher forcing. In contrast, our FSP-RevLM incorporates a reverse model with a different objective: mitigating dependence on teacher forcing. While trained on standard left-to-right sequences, FSP-RevLM aligns the forward model’s embeddings with those of a reverse model, effectively distilling the reverse order into the forward language model.
The closest prior work on reducing teacher forcing is Twin Networks (Serdyuk et al., 2017), which trains a reverse RNN to align forward hidden states to those of the reverse model to encourage long-range future dependence. While similar at a high level, our contributions go beyond the specific mechanism of generating future summaries via reverse language models. We cast future summary prediction as a broader framework for pretraining objective design, and provide evidence for when simple handcrafted summaries fail and when learned summaries from a reverse LM are needed. Moreover, extending this idea for large-scale pretraining with transformers is itself non-trivial. Just as Twin Networks anticipated aspects of framework, works predating modern MTP (Gloeckle et al., 2024) had already considered predicting multiple future tokens or parallel token blocks from a prefix (Tschannen et al., 2023; Monea et al., 2023), but did not establish such objectives as effective for large-scale transformer pretraining. Gloeckle et al. (2024) made that step for modern LLMs, and our work extends it by showing that FSP can be scaled for better pretraining.
Finally, Meet-in-the-Middle (MiM) (Nguyen et al., 2023) jointly trains forward and reverse models with shared parameters and aligns their output distributions as a regularizer. Our method differs in two key aspects: (1) we perform distillation with the reverse model rather than parameter sharing, and (2) we avoid the somewhat impractical inductive bias of MiM that the forward and reverse output distributions must match exactly.
@@ 6 Conclusion
In this work, we highlighted a key limitation of existing multi-token prediction methods: the difficulty of scaling auxiliary heads for long-horizon future prediction. Towards this, we proposed Future Summary Prediction (FSP), a novel pretraining framework that shifts the auxiliary objective from predicting specific future tokens to predicting a (learned) summary of the future. Experiments on 8B models, using both hand-crafted and learned summaries, demonstrate that FSP delivers a stronger, more robust training signal, improving over NTP and MTP on challenging reasoning and coding tasks. Our findings indicate that focusing on abstract, predictable aspects of the future is a promising strategy for designing more efficient and effective pretraining objectives for next-generation large language models. Future work can focus on designing computationally efficient approaches for learning future summaries (embeddings), potentially without training a reverse model.
@@ Acknowledgments
The authors thank Sachin Mehta for suggestions on the pretraining experiment setup, and Vaishnavh Nagarajan and Nanda H Krishna for their detailed feedback on the draft. The authors also gratefully acknowledge helpful discussions with Aniket Didolkar, Andrei Nicolicioiu, Mathurin Videau, Sarthak Mittal, Sharut Gupta, Vineet Jain, and Moksh Jain. This research was fully funded by Meta under the AI Mentorship (AIM) program with Mila. Ioannis Mitliagkas in his role as Divyat Mahajan’s academic advisor, acknowledges support by an NSERC Discovery grant (RGPIN-2019-06512), and Canada CIFAR AI chair. Divyat Mahajan acknowledges support via FRQNT doctoral scholarship (https://doi.org/10.69777/354785) for his graduate studies.